# Automatically Detecting Checked-In Secrets in Android Apps: How Far Are We?

### v1

@@ 4.1.3. Comparison of the Detection Capability of Each Tool
After manually analyzing the checked-in secrets of Google and Twitter, we discovered that for Google, Neither LeakScope or the Three-Layer Filter falsely detected checked-in secrets. Their precisions all reached 100%. The high precision of each tool is expected as the Google API key exhibits a strong pattern that can be precisely captured by regular expressions. In addition, both selected tools consider the reduction of false positives to be one of their main objectives. Three-Layer Filter leverages entropy filter, word filter and pattern filter to reduce false positives. LeakScope identifies checked-in secrets via backward slicing starting from the arguments of cloud APIs, ensuring that the extracted strings are checked-in secrets.
@@
However, we made different observations for Twitter client ID/Secret. The 10 Twitter checked-in secrets detected by LeakScope are all considered to be valid Twitter checked-in secrets through manual analysis. In contrast, none of the 383 Twitter checked-in secrets identified by the Three-Layer Filter were considered valid Twitter checked-in secrets. These checked-in secrets are indeed random strings, with some appearing to be tokens for services such as Facebook, CRASHLYTICS, or others. However, none of them are related to Twitter. Such high false positive rates occur because Twitter checked-in secrets are detected with loose regular expressions. This suggests that the effectiveness of the Three-Layer Filter heavily relies on the quality of the regular expressions. For the checked-in secrets that do not exhibit strong patterns, adopting the Three-Layer Filter is likely to result in a large number of false positives.
@@
Exploitable Checked-in Secrets. We acknowledge that our manual analysis may be subject to biases. Our manual analysis process may falsely consider a string to be a valid checked-in secret or vice versa. To address this concern, we validated the Google API keys by sending requests. We implement this validation process to demonstrate the exploitability of the detected checked-in secrets. After validating all the obtained Google API keys by sending a request to the server. 74 Google API keys grant access to the Google Map service. We reported these Google API keys to the developers of the corresponding Android apps, and we received responses from 13 Android apps that acknowledged the issue. This validation signifies that most of the checked-in secrets detected in this study are likely to be exploitable, revealing the severity of the checked-in secrets leakage in Android apps.
@@
4.4. RQ2.3-False Negative
4.4.1. RQ2.3-Methodology
In this research question, we aim to compare the false negatives of each tool. It is impractical to obtain the ground truth of our dataset as it requires manual labelling of over 1 million strings. In addition, all these strings are used in obfuscated code with minimal meaningful contexts. As a result, we cannot calculate the exact recall of each tool. To address this problem, we evaluate the overlap between the detection results of the LeakScope and the Three-Layer Filter. PassFinder was not included as it did not detect any checked-in secrets.
Furthermore, as described in the Section 4.1.3. We could not equalize the detection capability of LeakScope and the Three-Layer Filter. As a result, we focused solely on checked-in secrets categorized as Google API keys for a fair comparison since both of the tools detected a significant number of valid Google API keys. Twitter client ID and secrets are excluded because the Three-Layer Filter failed to detect any valid Twitter client and secrets after manual validation.
4.4.2. RQ2.3-Result
The Venn Diagram in Figure 1 shows the overlap of the detected checked-in secrets between each tool. The Three-Layer Filter identified 374 Google API keys, 11 of which were also detected by LeakScope. Conversely, LeakScope identified 1,816 Google API keys, with 11 overlapping with those found by the Three-Layer Filter. The minimal overlap between the tools suggests that both of them have missed detecting a significant number of checked-in secrets. Even with the absence of the ground truth, this signifies that all the tools have a relatively low recall. However, based on the overlap, we can still conclude that LeakScope missed fewer checked-in secrets compared to others.
@@
Figure 1. Venn Diagram for overlap of the number of unique detected Google API keys for LeakScope and Three-Layer Filter.
4.5. A Further Evaluation of PassFinder
@@ 4.5.1. Manual Inspection
In our experiment, the password model of PassFinder identified 11,343 strings as potential checked-in secrets. However, the context model concluded that none of these were valid checked-in secrets. To investigate further, we aimed to determine the validity of these strings and explore the reasons behind the context model’s failure to identify any of them. Therefore, we first randomly sampled 372 strings with their contextual surroundings, achieving a 95% confidence level with a 5% margin of error (ran, 2024) from the 11,343 strings identified by the password model. Upon manual investigation, we found that it is similar to the Three-Layer Filter with loose regular expressions; although these 372 strings are randomly generated strings, their context did not provide sufficient information to support that they are valid checked-in secrets. This was primarily due to obfuscation, which significantly differed from the training data used for the context models. The training set consisted of source code extracted from GitHub that contained meaningful, unobfuscated terms. Additionally, the extracted context was presented in Jimple format, an intermediate representation of Java, which further contributed to PassFinder’s ineffectiveness. To evaluate the impact of the Jimple format, we conducted an additional study where the Jimple format was transformed into the Java format. The details of the study are detailed in the following section.
4.5.2. Jimple to Java Transformation
@@
The use of the Jimple format instead of Java may have impacted the performance of the context model. Moreover, since the original sample did not include any valid checked-in secrets, it is challenging to fully determine the effectiveness of the context model. To address this, we created an additional test set containing 664 valid checked-in secrets from Google Cloud services detected by the Three-Layer Filter, along with their contextual surroundings manually converted into Java format. The conversion process was straightforward, as Jimple is an intermediate representation of Java. Listing 4.5.2 shows a Jimple-formatted code snippet from an Android app, closely resembling Java code and providing sufficient information to reconstruct the corresponding Java code. Listing 4.5.2 presents the transformed Java-format snippet from Listing 4.5.2.
{listing}
Extracted Code snippet in Jimple format
{minted}
[…]
4.5.3. Password Model on Labelled dataset
@@
We also conducted an additional experiment to fully validate the performance of the password model as the random sample strings in the previous section did not contain any valid checked-in secrets upon manual verification. Using the same test set from the previous section. This test aimed to evaluate the model’s ability to identify 664 valid checked-in secrets. However, the password model only identified 62 checked-in secrets from the test set, underscoring its restricted capability. The underperformance of the password model may be attributed to the nature of its training set. The machine-generated passwords in the training set were created using a random string generator (Feng et al., 2022), which likely differs from the structure and characteristics of actual checked-in secrets.
5. RQ3: Implications
@@
5.1. Limitations of the Evaluated Tools
@@
5.1.1. Limitation of Three Layer Filter (Intrinsic Value Analysis)
We identify three major limitations of the Three-Layer Filter approach.
@@ 5.1.1. Limitation of Three Layer Filter (Intrinsic Value Analysis)
First, the detection capability of the Three-Layer Filter is constrained by the manually crafted regular expressions. Regular expression matching is the most important step in the Three-Layer Filter. The number of manually crafted regular expressions determines the detection capability of the Three-Layer Filter. The strings identified by a precise regular expression will likely not require the following filtering process. In our experiment, only one string was excluded from the strings that were obtained through precise regular expressions. However, acquiring precise regular expressions for all categories of checked-in secret can be challenging, and some cloud providers may also generate checked-in secrets that do not exhibit a strong pattern (e.g., Twitter secrets). As a result, it is challenging to expand the detection capability of the Three-Layer Filter.
@@
Second, valid checked-in secrets can be excluded in the arbitrary filtering steps based on entropies, patterns and words. The entropy and pattern filter in our experiment excluded ten strings that are very likely to be a potential checked-in secret through our manual analysis. The threshold of the pattern sequence is set to be a length of four, which is problematic as some checked-in secret could contain such a pattern. For example, the checked-in secret of Firebase Messaging always starts with four repetitive “A”. While the word filter in our experiment did not produce any false negative cases, it’s plausible that checked-in secret could contain English words as one can never make an assumption about the randomly generated string. Furthermore, some checked-in secrets can be stored in a string with a format such as ”key=AIza…”. These checked-in secrets will be excluded as they did not exhibit any regular expression and contain English words such as ”key” and ”token”. As a result, it is hard to define accurate filters to eliminate the noises generated by the regular expressions without excluding false negatives.
@@
Finally, regular expressions cannot identify checked-in secrets that are computed from several strings. After analyzing the checked-in secrets detected exclusively by LeakScope, but which fall within the detection capability of the Three-Layer Filter (same category), we found that the Three-Layer Filter fails to extract checked-in secrets that are obtained through string manipulation methods such as append(). The presence of string manipulation poses a significant limitation for regular expression matching. A critical factor in regular expression matching is the range of the string length. If a checked-in secret is split in the Android apps, it would be impossible for the regular expression to match this string as the length is varied.
5.1.2. Limitation of LeakScope (Static analysis)
We identify four major limitations of LeakScope.
The primary limitation of LeakScope can only detect checked-in secrets from a list of pre-defined cloud APIs. To improve the detection capability of LeakScope, huge manual efforts are needed to expand the list of cloud API signatures. This largely limits the capability of LeakScope.
@@ 5.1.2. Limitation of LeakScope (Static analysis)
Additionally, since LeakScope uses Soot for backward analysis, it is also subject to the limitations of Soot. Other than the failures caused by the missing files during the reverse engineering process, LeakScope also failed to extract the exact value checked-in secrets that are stored as static variables within a Java class. e.g., public static String Google API key = "Alzaxxxx..". Soot sometimes can only get the name of these variables.
Third, the incorrect Cloud API signature generation and matching can impact the accuracy of detection.. We observed that LeakScope occasionally mismatches the Cloud API in the APK file. Some irrelevant APIs that have the same package structure and similar content would also be considered as one of the pre-defined cloud APIs during the analysis phase of LeakScope.
@@
5.1.3. Limitation of PassFinder (Machine-Learning-Based Technique)
When combining the predictions made by both the context model and the password model, PassFinder did not detect any checked-in secrets. To understand the reason behind this, we analyzed the two models separately. We observed that the poor performance of PassFinder can be attributed to (1) the training set and (2) the context modelling under obfuscation and insufficient information from 12 lines of context.
@@ 5.1.3. Limitation of PassFinder (Machine-Learning-Based Technique)
For the context model, even though we transformed code snippets from Jimple format to Java format within the training set, the effectiveness of these snippets was diminished by obfuscation. May add some pointers to the previous section where we discussed the further experiments on PassFinder This process replaced variable, method, and class names with random, meaningless letter sequences, reducing their overall significance. Furthermore, we observed that simply 12 lines of code could not always give enough information to detect the presence of checked-in secrets.
5.2. Implications: Future Direction
5.2.1. Complimentary among the tools
As described in Section 4.4, all the tools missed detecting a significant number of checked-in secrets, indicating that the actual recall of each tool is low. This suggests that future research should focus on improving the recall of checked-in secret detection techniques. Given the small overlap between the Three-Layer Filter and LeakScope, it is reasonable to assume that combining these two techniques could enhance the recall of checked-in secret detection.
However, these two techniques share a common limitation: it is difficult to expand their detection capability. The number of regular expressions constrains the Three-Layer Filter’s detection capability, while LeakScope’s detection capability is constrained by the list of pre-defined cloud API signatures. Both regular expressions and cloud APIs are challenging to discover and formalize.

### v2

@@ 4.1.3 Comparison of the Detection Capability of Each Tool
After manually analyzing a total of 2,117 Google API keys (1,816 detected by LeakScope and 664 by the Three-Layer Filter) and 393 Twitter secrets (10 from LeakScope and 383 from the Three-Layer Filter), we discovered that for Google, neither LeakScope nor the Three-Layer Filter falsely detected checked-in secrets. Their precision all reached 100%. In addition, both selected tools consider the reduction of false positives to be one of their main objectives. Three-Layer Filter leverages entropy filter, word filter, and pattern filter to reduce false positives. LeakScope identifies checked-in secrets via backward slicing starting from the arguments of cloud APIs, ensuring that the extracted strings are checked-in secrets.
@@
However, we made different observations for the Twitter client ID/Secret. The 10 Twitter checked-in secrets detected by LeakScope are all considered to be valid Twitter checked-in secrets through manual analysis. In contrast, none of the 383 Twitter checked-in secrets identified by the Three-Layer Filter were considered valid Twitter checked-in secrets. Upon closer inspection, these strings appeared random but generally fell into two categories: (1) 68% are tokens associated with other services such as Facebook or Crashlytics, and (2) 32% are benign random strings unrelated to authentication, such as session identifiers or serial numbers. This high false positive rate stems from the use of loose regular expressions for Twitter in the Three-Layer Filter, which capture any random string of a given length.
These findings highlight a core limitation of the Three-Layer Filter: while it can distinguish random from ordinary strings, randomness alone is insufficient to determine whether a string represents a valid secret. More broadly, this illustrates the weakness of methods that rely only on intrinsic string characteristics, as they tend to overgeneralize and misclassify benign data when consistent and recognizable patterns are absent.
@@
Exploitable Checked-in Secrets. We acknowledge that our manual analysis may be subject to biases. Our manual analysis process may falsely consider a string to be a valid checked-in secret or vice versa. To address this concern, we validated the Google API keys by sending requests. We implement this validation process to demonstrate the exploitability of the detected checked-in secrets. After validating all the obtained Google API keys by sending a request to the server, 74 Google API keys granted access to the Google Map service. We reported these Google API keys to the developers of the corresponding Android apps, and we received responses from 13 Android apps that acknowledged the issue. This validation signifies that most of the checked-in secrets detected in this study are likely to be exploitable, revealing the severity of the checked-in secrets leakage in Android apps.
@@
4.4 RQ2.3-False Negative
4.4.1 RQ2.3-Methodology
In this section, we aim to compare the false negatives of each tool. It is impractical to obtain the ground truth of our dataset as it requires manual labelling of over 1 million strings. In addition, all these strings are used in obfuscated code with minimal meaningful contexts. As a result, we cannot calculate the exact recall of each tool. To address this problem, we evaluate the overlap between the detection results of the LeakScope and the Three-Layer Filter. PassFinder was not included as it did not detect any checked-in secrets.
Furthermore, as described in the Section 4.1.3. We could not equalize the detection capability of LeakScope and the Three-Layer Filter. As a result, we focused solely on checked-in secrets categorized as Google API keys for a fair comparison, since both of the tools detected a significant number of valid Google API keys. Additionally, to ensure fairness, we removed duplicate Google API keys detected within the same tool, resulting in a final set of 312 unique keys for the Three-Layer Filter. Twitter client ID and secrets are excluded because the Three-Layer Filter failed to detect any valid Twitter client and secrets after manual validation.
4.4.2 RQ2.3-Result
The Venn Diagram in Figure 2 shows the overlap of the detected checked-in secrets between each tool. The Three-Layer Filter identified 374 Google API keys, 11 of which were also detected by LeakScope. Conversely, LeakScope identified 1,816 Google API keys, with 11 overlapping with those found by the Three-Layer Filter. The minimal overlap between the tools suggests that both of them have missed detecting a significant number of checked-in secrets. Even with the absence of the ground truth, this signifies that all the tools have a relatively low recall. However, based on the overlap, we can still conclude that LeakScope missed fewer checked-in secrets compared to the Three-Layer Filter.
@@
Figure 2: Venn Diagram for overlap of the number of unique detected Google API keys for LeakScope and Three-Layer Filter.
4.5 A Further Evaluation of PassFinder
@@ 4.5.1 Manual Inspection
In our experiment, the password model of PassFinder identified 11,343 strings as potential checked-in secrets. However, the context model concluded that none of these were valid checked-in secrets. To investigate further, we aimed to determine the validity of these strings and explore the reasons behind the context model’s failure to identify any of them. Therefore, we first randomly sampled 372 strings with their contextual surroundings, achieving a 95% confidence level with a 5% margin of error randomsample from the 11,343 strings identified by the password model. Upon manual investigation, we found that it is similar to the Three-Layer Filter with loose regular expressions; although these 372 strings are randomly generated strings, their context did not provide sufficient information to support that they are valid checked-in secrets. This was primarily due to obfuscation, which significantly differed from the training data used for the context models. The training set consisted of source code extracted from GitHub that contained meaningful, unobfuscated terms. Additionally, the extracted context was presented in Jimple format, an intermediate representation of Java, which further contributed to PassFinder’s ineffectiveness. To evaluate the impact of the Jimple format, we conducted an additional study where the Jimple format was transformed into the Java format. The details of the study are detailed in the following section.
4.5.2 Jimple to Java Transformation
@@
The use of the Jimple format instead of Java may have impacted the performance of the context model. Moreover, since the original sample did not include any valid checked-in secrets, it is challenging to fully determine the effectiveness of the context model. To address this, we created an additional test set containing 664 valid checked-in secrets from Google Cloud services detected by the Three-Layer Filter, along with their contextual surroundings manually converted into Java format. The conversion process was straightforward, as Jimple is an intermediate representation of Java. Listing 3 shows a Jimple-formatted code snippet from an Android app, closely resembling Java code and providing sufficient information to reconstruct the corresponding Java code. Listing 4 presents the transformed Java-format snippet from Listing 3.
Listing 3: Extracted code snippet in Jimple format
1\$r0 = newarray (java.lang.String)[3]
2\$r0[0] = ""
3\$r0[1] = "67"
4\$r0[2] = ""
5<jp.b: java.lang.String[] c> = \$r0
Listing 4: Transformed Java code snippet from Listing 3
1String[] array = new String[3];
2array[0] = "";
3array[1] = "67";
4array[2] = "";
5jp.b.c = array;
The context model still did not identify any from the test set. After analyzing the Java code snippets from the training set, we found that the majority of the code snippets that contain passwords are the declaration of global variables of classes. In contrast, the code snippets containing the checked-in secrets in APK files are the locations where the checked-in secrets are used, such as the example we discuss in Listing 1. This indicates a fundamental difference between the contexts in the training set and those in the test set. Specifically, the training data does not adequately represent the usage patterns of checked-in secrets found in APK files. This misalignment may explain the context model’s inability to detect these secrets effectively.
4.5.3 Password Model on Labelled dataset
@@
We also conducted an additional experiment to fully validate the performance of the password model, as the random sample strings in the previous section did not contain any valid checked-in secrets upon manual verification. Using the same test set from the previous section. This test aimed to evaluate the model’s ability to identify 664 valid checked-in secrets. However, the password model only identified 62 checked-in secrets from the test set, underscoring its restricted capability. The underperformance of the password model may be attributed to the nature of its training set. The machine-generated passwords in the training set were created using a random string generator passfinder, which likely differs from the structure and characteristics of actual checked-in secrets.
5 RQ3: Implications
@@
5.1 Limitations of the Evaluated Tools
@@
5.1.1 Limitation of Three Layer Filter (Intrinsic Value Analysis)
@@
First, the detection capability of the Three-Layer Filter is constrained by the manually crafted regular expressions. Regular expression matching is the most important step in the Three-Layer Filter. The number of manually crafted regular expressions determines the detection capability of the Three-Layer Filter. The strings identified by a precise regular expression will likely not require the following filtering process. In our experiment, only one string was excluded from the strings that were obtained through precise regular expressions. However, acquiring precise regular expressions for all categories of checked-in secret can be challenging. Some secrets, such as Google API keys, tend to follow recognizable and consistent formats, making them easier to capture with well-defined patterns. In contrast, secrets issued by other providers, such as Twitter Client IDs or Secrets, often consist of arbitrary alphanumeric strings without a stable structure. For instance, Twitter secrets may appear as random strings like “xvz1******Bog” or “L8qq9PZ******0vZk”, which are difficult to distinguish from non-secret strings using regex alone. As a result, it is challenging to expand the detection capability of the Three-Layer Filter.
@@
Second, valid checked-in secrets can be excluded in the arbitrary filtering steps based on entropies, patterns, and words. The entropy and pattern filter in our experiment excluded ten strings that are very likely to be a potential checked-in secret through our manual analysis. The threshold of the pattern sequence is set to be a length of four, which is problematic as some checked-in secret could contain such a pattern. A concrete example is Firebase Cloud Messaging API keys, which often begin with four identical characters such as “AAAA…”. These legitimate keys were mistakenly excluded due to the repetitive pattern.
Similarly, while the word filter did not lead to any false negatives in our dataset, it poses a risk of filtering out valid keys that happen to include common English words. For instance, a checked-in secret stored in a string like “key=AIzaSyB…”or “token=abc123…” could be excluded simply because it includes substrings like “key” or “token”, even though the value is a valid checked-in secret. Since randomly generated keys can occasionally contain English words by chance, it’s difficult to establish filtering rules that are both strict enough to eliminate noise and flexible enough to avoid excluding true positives. As a result, designing precise and reliable filters remains a significant challenge.
@@
Finally, regular expressions cannot identify checked-in secrets that are computed from several strings. After analyzing the checked-in secrets detected exclusively by LeakScope, but which fall within the detection capability of the Three-Layer Filter (same category), we found that the Three-Layer Filter fails to extract checked-in secrets that are obtained through string manipulation methods such as append(). For example, if an checked-in secret is constructed in an Android app using multiple string concatenation operations like “AIza” appended with “SyA…”, the resulting key will not exist in its entirety in the static string pool, making it invisible to regex-based detection. Regular expressions also rely heavily on the total length of the string to enforce structure, and when keys are split or dynamically assembled at runtime, their fragmented form falls outside the expected pattern length, preventing a successful match. This limitation underscores the need for detection techniques that can handle non-literal and dynamically constructed secrets.
5.1.2 Limitation of LeakScope (Static Analysis)
The primary limitation of LeakScope can only detect checked-in secrets from a list of pre-defined cloud APIs. To improve the detection capability of LeakScope, significant manual effort is required to expand and maintain the list of cloud API signatures. This dependency on predefined APIs substantially limits its ability to generalize or detect secrets outside the curated set. For example, during our study, we attempted to identify any cloud APIs that accept PayPal checked-in secrets as input but were unable to find a relevant API call to include in the signature list. As a result, even if a PayPal API key were present in the code, LeakScope would be unable to detect it due to the absence of a matching API in its predefined list. This example illustrates the inherent limitation of relying on a fixed set of API signatures for detection.
@@ 5.1.2 Limitation of LeakScope (Static Analysis)
Additionally, since LeakScope uses Soot for backward analysis, it is also subject to the limitations of Soot. Other than the failures caused by the missing files during the reverse engineering process, LeakScope also failed to extract the exact value checked-in secrets that are stored as static variables within a Java class. e.g., public static String Google API key = “Alza****..”. Soot sometimes can only get the name of these variables.
Third, the incorrect Cloud API signature generation and matching can impact the accuracy of detection. We observed that LeakScope occasionally mismatches the Cloud API in the APK file. Some irrelevant APIs that have the same package structure and similar content would also be considered as one of the pre-defined cloud APIs during the analysis phase of LeakScope. This limitation arises because LeakScope primarily relies on package structures when generating cloud API signatures. Consequently, methods from unrelated libraries that share similar package hierarchies or naming conventions may be misclassified as valid API calls. For example, we found that the method setOAuthConsumerKey(), which is typically associated with Twitter API usage, was sometimes incorrectly matched to unrelated methods residing in third-party libraries that resemble the structure of the official Twitter SDK. This kind of false matching can lead to inaccurate detection results and highlights the challenge of relying solely on static API signature matching without additional semantic validation.
@@
5.1.3 Limitation of PassFinder (Machine-Learning-Based Technique)
When combining the predictions made by both the context model and the password model, PassFinder did not detect any checked-in secrets. To understand the reason behind this, we analyzed the two models separately.
We observed that the poor performance of PassFinder can be attributed to (1) the training set and (2) the context modelling under obfuscation and insufficient information from 12 lines of context.
@@ 5.1.3 Limitation of PassFinder (Machine-Learning-Based Technique)
For the context model, even though we transformed code snippets from Jimple format to Java format within the training set, the effectiveness of these snippets was diminished by obfuscation. This process replaced variable, method, and class names with random, meaningless letter sequences, reducing their overall significance. Furthermore, we observed that simply 12 lines of code could not always give enough information to detect the presence of checked-in secrets.
Summary: In summary, each analyzed method exhibits intrinsic limitations that are not specific to Android but are fundamental to their underlying detection methodologies. However, the complexity of Android APK structures and the pervasive use of obfuscation introduce additional challenges that further constrain their performance. The Three-Layer Filter is limited by its reliance on manually defined regular expression rules and its assumption that secrets appear as complete string literals in source code. These constraints prevent it from recognizing secrets that are dynamically constructed, encoded, or transformed at runtime. In the Android environment, the effectiveness of this approach can be further reduced by obfuscation or string encoding, which alter literal values or conceal them from direct inspection. In some cases, missing or corrupted files during decompilation may also lead to incomplete coverage of code units, further limiting the Three-Layer Filter’s detection scope. LeakScope, while more sophisticated in design, is strongly affected by the characteristics of the Android environment. Its reliance on static analysis and predefined API signatures restricts detection to known APIs and simple invocation structures. In practice, the event-driven architecture of Android applications, combined with the complexity of component interactions and the widespread use of code obfuscation further hinder data-flow propagation. For instance, among 3,179 identified cloud API usages, only 1,846 corresponding values were successfully extracted. Additionally, we observed five cases where incorrect APIs were used for further data flow analysis due to mismatches caused by code obfuscation. Quantitatively, the overlap between the Three-Layer Filter and LeakScope detections among all identified Google API keys is only 0.5% as shown in Figure 2, indicating that these deterministic methods capture largely distinct subsets of secrets and that many valid cases remain undetected because of their differing assumptions. Machine-learning-based approaches, such as PassFinder, provide greater flexibility by learning diverse representations of secret usage rather than relying on predefined patterns. Nevertheless, pervasive obfuscation in production Android apps significantly degrades their performance by altering or removing contextual cues essential for effective learning. These challenges call the need for Android-aware learning strategies, which motivated our preliminary work on String Group Analysis, a framework designed to enhance the robustness of learning-based detection under obfuscation.
5.2 Future Direction
5.2.1 Complimentary Results of the Tools
As described in Section 4.4, all the tools missed detecting a significant number of checked-in secrets, indicating that the actual recall of each tool is low. The minimal overlap between the Three-Layer Filter and LeakScope stems from their fundamentally different detection strategies. The Three-Layer Filter targets complete strings stored directly in the code, making it effective for identifying hardcoded, static secrets. In contrast, LeakScope applies inter-procedural analysis with control and data-flow reasoning, which enables it to capture secrets that are dynamically constructed or spread across program elements. Consequently, LeakScope can recover a broader range of string values.
Despite these strengths, both approaches have inherent limitations that restrict their detection capabilities. The Three-Layer Filter is constrained by the coverage of its regular expressions, while LeakScope relies on a predefined set of cloud API signatures. While effective for cloud-related secrets, this reliance risks overlooking other formats, such as secrets embedded in custom requests or transmitted through non-standard APIs. Moreover, both regular expressions and cloud API specifications are difficult to comprehensively enumerate and formalize.
These trade-offs highlight the complementary nature of the two techniques. LeakScope’s data-flow analysis can help reconstruct or complete string values that are otherwise inaccessible to regex-based matching, thereby enhancing the effectiveness of the Three-Layer Filter. At the same time, regex-based approaches can efficiently detect fully formed strings that may not be directly associated with known API usage. Together, such methods can offer improved recall by mitigating each other’s blind spots.
This complementary relationship is also reflected in the quantitative results. Although calculating exact false-negative rates is infeasible due to the absence of a ground-truth benchmark for checked-in secrets in Android apps, additional quantitative analysis was conducted to better illustrate the complementarity between LeakScope and the Three-Layer Filter. Among all valid Google API keys detected, only 11 were found by both tools as shown in Figure 2, which is a minimal overlap account for 0.5% of all detected Google API keys. This indicates that they capture largely distinct subsets of secrets. In particular, 99.4% of LeakScope’s detections and 97.1% of the Three-Layer Filter’s detections are non-overlapping, and combining the two approaches increases overall detection coverage by roughly 20% compared with using LeakScope alone. These quantitative results reinforce that integrating syntactic- and flow-based analyses yields broader coverage and complementary detection capabilities.